# EVENet: Evidence-based Ensemble Learning for Uncertainty-aware Brain Parcellation Using Diffusion MRI

### v1

@@ Abstract
In this study, we developed an Evidence-based Ensemble Neural Network, namely EVENet, for anatomical brain parcellation using diffusion MRI. The key innovation of EVENet is the design of an evidential deep learning framework to quantify predictive uncertainty at each voxel during a single inference. Using EVENet, we obtained accurate parcellation and uncertainty estimates across different datasets from healthy and clinical populations and with different imaging acquisitions. The overall network includes five parallel subnetworks, where each is dedicated to learning the FreeSurfer parcellation for a certain diffusion MRI parameter. An evidence-based ensemble methodology is then proposed to fuse the individual outputs. We perform experimental evaluations on large-scale datasets from multiple imaging sources, including high-quality diffusion MRI data from healthy adults and clinically diffusion MRI data from participants with various brain diseases (schizophrenia, bipolar disorder, attention-deficit/hyperactivity disorder, Parkinson’s disease, cerebral small vessel disease, and neurosurgical patients with brain tumors). Compared to several state-of-the-art methods, our experimental results demonstrate highly improved parcellation accuracy across the multiple testing datasets despite the differences in dMRI acquisition protocols and health conditions. Furthermore, thanks to the uncertainty estimation, our EVENet approach demonstrates a good ability to detect abnormal brain regions in patients with lesions, enhancing the interpretability and reliability of the segmentation results.
@@ 1 Introduction
Advanced techniques have been developed for direct brain parcellation from dMRI data using deep learning, eliminating the need for inter-modality registration and thus enhancing parcellation performance (Liu et al., 2007; Wen et al., 2013; Yap et al., 2015; Ciritsis et al., 2018; Zhang et al., 2021; Theaud et al., 2022; Zhang et al., 2023), with improved accuracy and efficiency (Theaud et al., 2022; Zhang et al., 2023; Billot et al., 2023). However, an important limitation of these advanced methods is that they were trained and tested on data that are closely matched in terms of subject characteristics and imaging acquisition protocols. Their effectiveness diminishes when applied to out-of-distribution data, i.e., data that is different from the training data. As a result, the generalizability of the existing dMRI parcellation methods to data from different sources remains a challenge. In general, there are several major sources of variability that lead to uncertainty and poorer generalizability in deep learning model predictions in dMRI parcellation. First, it is well known that variations across acquisition sites, arising from different MRI scanners and/or acquisition protocols can result in significant biases in dMRI data (Billot et al., 2023; Mirzaalian et al., 2016; Karayumak et al., 2019; Cetin Karayumak et al., 2018; Ning et al., 2020; Tax et al., 2019; Li et al., 2024). Site variability can result in poor generalizability to data from different centers/sites/scanners. Second, the parcellation performance can be compounded by differences in spatial resolution, leading to varying overlaps or partial volumes between anatomical structures (Guevara et al., 2020). Consequently, the certainty in the parcellation prediction in the boundary between regions or tissue types diminishes. Third, current parcellation methods are primarily designed and/or trained on the brains of healthy individuals with normal structural appearances. As a result, they can be erroneous when applied to brains with abnormalities such as lesions, brain tumors, and white matter hyperintensity (WMH). Parcellation of out-of-distribution data (e.g., Figure 1 with a dataset containing a glioma) is challenging not just for deep learning approaches but also for traditional parcellation tools such as FreeSurfer (Fischl, 2012).
@@
Evidential deep learning (Sensoy et al., 2018; Zou et al., 2022; Li et al., 2023), rooted in the principles of subjective logic theory (Jsang, 2018), offers a simple yet effective tool for uncertainty estimation. Importantly, this relatively new approach provides improved results, especially for the classification of out-of-distribution samples (Sensoy et al., 2018). So far, evidential deep learning has been used for few-class segmentation such as binary tumor segmentation (Zou et al., 2022; Li et al., 2023), but holds great potential for multi-class segmentation (e.g. parcellation of the brain image into over tens of regions) for its high efficiency and ease of implementation. Subjective logic theory extends traditional probability theory by introducing the notion of uncertainty that integrates evidence and belief (Jsang, 2018). Unlike standard logic, where propositions are dichotomously considered true or false, or probabilistic logic, which assigns probabilities in the range [0, 1] to all arguments, evidential deep learning explicitly acknowledges a probability for uncertainty (Sensoy et al., 2018). In this framework, the network is trained to simultaneously optimize data fit while maximizing evidence, hence minimizing uncertainty.
@@
In this study, we propose the Evidence-based Ensemble Neural Network (EVENet), a novel uncertainty-aware deep learning method for anatomical brain parcellation of cortical and subcortical regions directly from dMRI data. The key innovation of EVENet is its utilization of evidential deep learning to quantify uncertainty at each voxel during a single inference. As a result, even though EVENet is trained using data from healthy individuals scanned with a single MRI protocol, it is able to provide accurate parcellation across different datasets with different imaging acquisitions, and from different populations, including individuals with brain disorders. In addition, we combine the uncertainty estimation across different channels, this way EVENet leverages multiple dMRI parameters to perform fine-scale parcellation, followed by an evidence-based ensemble fusion of the output from each dMRI parameter for final parcellation predictions and uncertainty estimation. Given that different diffusion parameters describe tissue properties from varying perspectives, incorporating uncertainty estimation can enhance prediction results by accounting for the differences in outputs from each parameter. We perform experimental evaluations on large-scale datasets from multiple imaging sources, including high-quality dMRI images from healthy adults and clinically dMRI data from participants with various brain disorders (including schizophrenia, bipolar disorder, attention-deficit/hyperactivity disorder, Parkinson’s disease, cerebral small vessel disease, and neurosurgical patients with brain tumors). Our results indicate not only improved parcellation accuracy but also the ability to detect abnormal brain regions, suggesting EVENet’s robustness and generalization capabilities in both research and clinical settings.
@@ 2.3 Evidence-based Ensemble Learning for Uncertainty Estimation
Leveraging the outputs obtained from the five sub-networks, the entropy based on the averaged beliefs is used for the final uncertainty estimation. The estimation is described as:
@@
where
$p_{m}^{{}^{\prime}c}=\frac{\sum_{q=1}^{M}p_{m}^{q,c}}{M}=\frac{\sum_{q=1}^{M}e_{m}^{q,c}}{M\times S_{m}}$ (9)
$M$ is the total number of subnetworks, and $N$ is the total number of class labels.
@@ 2.4 Implementation and Parameter Settings
Our method is implemented using Pytorch 2.0 (Paszke et al., 2019) and trained on a server equipped with NVIDIA RTX3090 GPUs. For each subnetwork, the development of our model is based on FastSurferCNN (https://github.com/Deep-MI/FastSurfer) and DDparcel (https://github.com/zhangfanmark/DDParcel). Code pieces in (Zou et al., 2022) are also used for the calculation of evidential loss (https://github.com/Cocofeat/TBraTS). We adopted the configurations such as learning rate and number of input channels as recommended by the FastSurferCNN paper, considering these values as a starting point for our model. We employed Adam as the optimizer (Kingma, 2014) because of its effectiveness in handling sparse gradients on noisy problems. A learning rate decay strategy was implemented to enhance training dynamics, with an initial rate of 0.01, then decreased by 95% every five epochs over 200 epochs. The models were iteratively updated using a batch size of 8, which balances the trade-off between memory constraints and the benefits of mini-batch training. For the overall loss, the parameters and kl are empirically set to 0.7 and 0.4, respectively, to balance the model’s focus between parcellation performance and quality of uncertainty estimation.
@@ 3.2.1 Ablation Study
First, the effectiveness of the ensemble criteria and backbone network is evaluated in an ablation study, where a comparison among the following methods is performed. The probability-based method averages prediction probabilities derived from the original softmax activation layers. The entropy-based method calculates the entropy of the output probabilities and selects the label with minimal entropy as the final prediction for each voxel. The evidence-based method (proposed) employs the evidence learning module and conducts ensemble segmentation as described in Section 2.3. For the former two methods, we use the default loss settings recommended in the original papers to train the subnetworks. For evidence-based method, we also incorporate evidential deep learning loss. We test these three criteria of the ensemble on three backbone networks, including FastSurfer, nnU-Net, and Swin UNETR. The evaluation is performed using the HCP testing data. For each compared method and each HCP testing subject, the Dice score, recall, and intersection over union (IoU) for each FS region is computed between the prediction and the ground truth parcellation, and then the average score across all regions is obtained.
@@ 5 Discussion
Potential limitations of the present study, including suggested future work to address limitations, are as follows. First, it is interesting to integrate other imaging modalities, such as T1-weighted and functional MRI, for a more comprehensive understanding of brain structure and function. Second, optimizing EVENet for real-time processing could benefit surgical planning and intraoperative guidance. Incorporating advanced neural network training techniques, such as federated learning or transfer learning, could enhance training efficiency and effectiveness, facilitating the use of decentralized data sources while respecting privacy concerns or adapting the network to new tasks with minimal retraining. Finally, our evaluation was also limited to a certain age range and did not extend to very young ages (e.g., babies and neonates). Future work should include the curation of training data that reflect the anatomy of more specific populations.
@@ 6 Conclusion
The data used in this project include the public HCP (www.humanconnectome.org), CNP (https://openfmri.org/dataset/ds000030), and PPMI (http://www.ppmi-info.org) datasets. The WMH and BT datasets are our in-house datasets, which are not publicly available because public availability would compromise participant confidentiality and participant privacy. The code and trained model will be made publically available at: https://github.com/chenjun-li/EVENet.

### v2

# DDEvENet: Evidence-based Ensemble Learning for Uncertainty-aware Brain Parcellation Using Diffusion MRI

@@ Abstract
In this study, we developed an Evidence-based Ensemble Neural Network, namely EVENet, for anatomical brain parcellation using diffusion MRI. The key innovation of EVENet is the design of an evidential deep learning framework to quantify predictive uncertainty at each voxel during a single inference. To do so, we design an evidence-based ensemble learning framework for uncertainty-aware parcellation to leverage the multiple dMRI parameters derived from diffusion MRI. Using EVENet, we obtained accurate parcellation and uncertainty estimates across different datasets from healthy and clinical populations and with different imaging acquisitions. The overall network includes five parallel subnetworks, where each is dedicated to learning the FreeSurfer parcellation for a certain diffusion MRI parameter. An evidence-based ensemble methodology is then proposed to fuse the individual outputs. We perform experimental evaluations on large-scale datasets from multiple imaging sources, including high-quality diffusion MRI data from healthy adults and clinically diffusion MRI data from participants with various brain diseases (schizophrenia, bipolar disorder, attention-deficit/hyperactivity disorder, Parkinson’s disease, cerebral small vessel disease, and neurosurgical patients with brain tumors). Compared to several state-of-the-art methods, our experimental results demonstrate highly improved parcellation accuracy across the multiple testing datasets despite the differences in dMRI acquisition protocols and health conditions. Furthermore, thanks to the uncertainty estimation, our EVENet approach demonstrates a good ability to detect abnormal brain regions in patients with lesions, enhancing the interpretability and reliability of the segmentation results.
@@ 1 Introduction
Advanced techniques have been developed for direct brain parcellation from dMRI data using deep learning, eliminating the need for inter-modality registration and thus enhancing parcellation performance (Liu et al., 2007; Wen et al., 2013; Yap et al., 2015; Ciritsis et al., 2018; Zhang et al., 2021; Theaud et al., 2022; Zhang et al., 2023), with improved accuracy and efficiency (Theaud et al., 2022; Zhang et al., 2023; Billot et al., 2023). Early work using deep learning typically employs encoder-decoder architecture with convolutional neural networks (CNNs) for efficient feature extraction and segmentation (Zhang et al., 2021; Theaud et al., 2022). More recent methods (Li et al., 2023a; Hayat and Aramvith, 2024) have considered the use of data fusion techniques with attention mechanisms to leverage more contextual information and improve accuracy. However, an important limitation of these advanced methods is that they were trained and tested on data that are closely matched in terms of subject characteristics and imaging acquisition protocols. Their effectiveness diminishes when applied to out-of-distribution data, i.e., data that is different from the training data. As a result, the generalizability of the existing dMRI parcellation methods to data from different sources remains a challenge. In general, there are several major sources of variability that lead to uncertainty and poorer generalizability in deep learning model predictions in dMRI parcellation. First, it is well known that variations across acquisition sites, arising from different MRI scanners and/or acquisition protocols can result in significant biases in dMRI data (Karayumak et al., 2019; Tax et al., 2019; Li et al., 2024).. Site variability can result in poor generalizability to data from different centers/sites/scanners. Second, the parcellation performance can be compounded by differences in spatial resolution, leading to varying overlaps or partial volumes between anatomical structures (Guevara et al., 2020). Consequently, the certainty in the parcellation prediction in the boundary between regions or tissue types diminishes. Third, current parcellation methods are primarily designed and/or trained on the brains of healthy individuals with normal structural appearances. As a result, they can be erroneous when applied to brains with abnormalities such as lesions, brain tumors, and white matter hyperintensity (WMH). Parcellation of out-of-distribution data (e.g., Figure 1 with a dataset containing a glioma) is challenging not just for deep learning approaches but also for traditional parcellation tools such as FreeSurfer (Fischl, 2012).
@@
Evidential deep learning (Sensoy et al., 2018; Zou et al., 2022; Li et al., 2023b), rooted in the principles of subjective logic theory (Jsang, 2018), offers a simple yet effective tool for uncertainty estimation. Importantly, this relatively new approach provides improved results, especially for the classification of out-of-distribution samples (Sensoy et al., 2018). So far, evidential deep learning has been used for few-class segmentation such as binary tumor segmentation (Zou et al., 2022; Li et al., 2023b), but holds great potential for multi-class segmentation (e.g. parcellation of the brain image into over tens of regions) for its high efficiency and ease of implementation. Subjective logic theory extends traditional probability theory by introducing the notion of uncertainty that integrates evidence and belief (Jsang, 2018). Unlike standard logic, where propositions are dichotomously considered true or false, or probabilistic logic, which assigns probabilities in the range [0, 1] to all arguments, evidential deep learning explicitly acknowledges a probability for uncertainty (Sensoy et al., 2018). In this framework, the network is trained to simultaneously optimize data fit while maximizing evidence, hence minimizing uncertainty.
@@
In this study, we propose the Evidence-based Ensemble Neural Network (EVENet), a novel uncertainty-aware deep learning method for anatomical brain parcellation of cortical and subcortical regions directly from dMRI data. The novelty of our approach lies primarily in the application of evidential deep learning combined with ensemble methods for uncertainty-aware brain parcellation. Our method is built on established neural network architectures and techniques but tailored to consider the multi-parameter property of dMRI data, enhancing both the accuracy and interpretability of parcellation results. The key innovation of EVENet is its utilization of evidential deep learning to quantify uncertainty at each voxel during a single inference. As a result, even though EVENet is trained using data from healthy individuals scanned with a single MRI protocol, it is able to provide accurate parcellation across different datasets with different imaging acquisitions, and from different populations, including individuals with brain disorders. In addition, we combine the uncertainty estimation across different channels, this way EVENet leverages multiple dMRI parameters to perform fine-scale parcellation, followed by an evidence-based ensemble fusion of the output from each dMRI parameter for final parcellation predictions and uncertainty estimation. Given that different diffusion parameters describe tissue properties from varying perspectives, incorporating uncertainty estimation can enhance prediction results by accounting for the differences in outputs from each parameter. We perform experimental evaluations on large-scale datasets from multiple imaging sources, including high-quality dMRI images from healthy adults and clinically dMRI data from participants with various brain disorders (including schizophrenia, bipolar disorder, attention-deficit/hyperactivity disorder, Parkinson’s disease, cerebral small vessel disease, and neurosurgical patients with brain tumors). Our results indicate not only improved parcellation accuracy but also the ability to detect abnormal brain regions, suggesting EVENet’s robustness and generalization capabilities in both research and clinical settings.
@@ 2.3 Evidence-based Ensemble Learning for Uncertainty Estimation
Leveraging the outputs obtained from the five sub-networks, the entropy based on the averaged beliefs is used for the final uncertainty estimation. Specifically, each subnetwork produces a set of beliefs $p_{m}^{q,c}$ and an associated uncertainty $u_{m}^{q}$ for each voxel $m$, where $q$ indexes the subnetwork corresponding to each dMRI parameter, and $c$ indexes the class labels. To combine the information from the multiple subnetworks, we first compute an average belief across all subnetworks for each class at each voxel:
$p_{m}^{{}^{\prime}c}=\frac{1}{M}\sum_{q=1}^{M}p_{m}^{q,c}$ (8)
where $N$ is the total number of class labels. The entropy $u^{\prime}_{m}$ quantifies the uncertainty in the final prediction at voxel $m$, with higher entropy indicating higher uncertainty. Next, we compute the entropy of the averaged belief distribution at each voxel to obtain the final uncertainty estimation:
@@
where $N$ is the total number of class labels. The entropy $u^{\prime}_{m}$ quantifies the uncertainty in the final prediction at voxel $m$, with higher entropy indicating higher uncertainty.
@@
This means that for each voxel, we rely on the subnetwork that is most confident (i.e., has the lowest uncertainty) in its prediction at that location. The rationale behind this choice is that different dMRI parameters may provide more reliable information in different regions due to their sensitivity to different microstructural properties. By selecting the prediction from the subnetwork with the lowest uncertainty, we aim to leverage the most trustworthy information available at each voxel.
Entropy-based uncertainty affects the final segmentation by providing a measure of confidence in the sub-network predictions, which can be used to inform further processing or decision-making. For instance, regions with high entropy may correspond to areas where the model is uncertain due to factors such as image noise, anatomical variability, or pathology. By identifying these regions or modalities, clinicians or researchers can focus their attention on areas that may require manual review or additional imaging.
@@ 2.4 Implementation and Parameter Settings
Our method is implemented using Pytorch 2.0 (Paszke et al., 2019) and trained on a server equipped with NVIDIA RTX3090 GPUs. For each subnetwork, the development of our model is based on FastSurferCNN (https://github.com/Deep-MI/FastSurfer) and DDparcel (https://github.com/zhangfanmark/DDParcel). Code pieces in (Zou et al., 2022) are also used for the calculation of evidential loss (https://github.com/Cocofeat/TBraTS). We adopted the configurations such as learning rate and number of input channels as recommended by the FastSurferCNN paper, considering these values as a starting point for our model. We employed Adam as the optimizer (Kingma, 2014) because of its effectiveness in handling sparse gradients on noisy problems. A learning rate decay strategy was implemented to enhance training dynamics, with an initial rate of 0.01, then decreased by 95% every five epochs over 200 epochs. The models were iteratively updated using a batch size of 8, which balances the trade-off between memory constraints and the benefits of mini-batch training. For the overall loss, the parameters and kl are empirically set to 0.7 and 0.4, respectively, to balance the model’s focus between parcellation performance and quality of uncertainty estimation. To improve the generalization, we applied data augmentation techniques during training. Specifically, we performed random rotations (±15 degrees), scaling (90% to 110%), and horizontal flipping on the input images. These augmentations were applied on-the-fly during training. The use of data augmentation resulted in a 2% increase in the Dice coefficient on the validation set, indicating enhanced model robustness.
@@ 3.2.1 Ablation Study
First, the effectiveness of the ensemble criteria and backbone network is evaluated in an ablation study, a comparison was performed among the following methods, including: (1) the probability-based method, (2) the entropy-based method, and (3) the evidence-based method (proposed). Specifically, the probability-based method averages prediction probabilities derived from the original softmax activation layers. The entropy-based method calculates the entropy of the output probabilities and selects the label with minimal entropy as the final prediction for each voxel. The evidence-based method (proposed) employs the evidence learning module and conducts ensemble segmentation as described in Section 2.3. For the former two methods, we use the default loss settings recommended in the original papers to train the subnetworks. For evidence-based method, we also incorporate evidential deep learning loss. We test these three criteria of the ensemble on three backbone networks, including FastSurfer, nnU-Net, and Swin UNETR. The evaluation is performed using the HCP testing data. For each compared method and each HCP testing subject, the Dice score, recall, and intersection over union (IoU) for each FS region is computed between the prediction and the ground truth parcellation, and then the average score across all regions is obtained.
@@ 5 Discussion
Potential limitations of the present study, including suggested future work to address limitations, are as follows. First, an interesting direction for future work is to integrate other imaging modalities such as T1-weighted and functional MRI into our framework. These modalities provide complementary structural and functional information that can further enhance the accuracy and robustness of uncertainty estimation. Multi-modal fusion strategies could be developed to leverage the strengths of each modality for a more comprehensive understanding of brain structure and function. Second, while our approach already offers a large computational advantage over the traditional registration-based approach, we acknowledge the potential for further optimization, particularly in the context of real-time processing for applications such as surgical planning and intraoperative guidance. Techniques such as model compression and pruning are promising directions to explore in the future, as they can reduce computational complexity. Although our current method meets the immediate demands of large-scale studies, further refinements in this area are feasible and will be considered in future work to enhance real-time capabilities. Third, recent advancements in self-supervised learning (Gui et al., 2024; Huang et al., 2023) and neural architecture search (NAS) (Kang et al., 2023; Qin et al., 2023) have shown promise in medical imaging applications. Self-supervised learning could enable EVENet to leverage unlabeled data for pre-training, potentially improving performance on limited datasets. NAS could help in discovering more efficient network architectures tailored to our specific task, possibly reducing computational overhead. Incorporating these approaches in future work may enhance both the efficiency and generalizability of EVENet. Finally, we tested our method on populations with a wide age range including children, young adults, and elderly adults (see Supplementary Figure 1). However, our evaluation was also limited to a certain age range and did not extend to very young ages (e.g., babies and neonates). It is well known that the myelination of white matter in neonates is essentially different from the populations under study. Therefore, the curation of training data that reflects the anatomy of the specific populations might be needed to further improve the generalizability of EVENet across different age groups and developmental stages.
@@ 6 Conclusion
The data used in this project include the public HCP (www.humanconnectome.org), CNP (https://openfmri.org/dataset/ds000030), and PPMI (http://www.ppmi-info.org) datasets. The raw imaging data of the WMH and BT datasets are not publicly available because public availability would compromise participant confidentiality and participant privacy, but the derived diffusion MRI parameter maps will be made available upon request. The code and trained model will be made publically available at: https://github.com/chenjun-li/EVENet.